one number exceeds another by 24. the sum of the numbers is 58. what are the numbers
Answer by checkley77(12569)

X-Y=24
X+Y=58
--------
2X=82
X=82/2
X=41 ANS.
41+Y=58
Y=58-41
Y=17 ANS.